OpenStudy (anonymous):

Write an equation for the line that passes through the points (0, -2) and (-4, 0). y= -2x- 4 y= -2x- 2 y= -1/2x - 2 y= -1/2x - 4

OpenStudy (anonymous):

@iheartfood is that the same??

OpenStudy (ivettef365):

The equation of the line is y=mx+b

OpenStudy (anonymous):

y= -(1/2)x-2

OpenStudy (anonymous):

point slope form:\[(y-y_1)=m(x-x_1)\] formula for slope:\[m={{y_2-y_1}\over{x_2-x_1}}\] equation of a line:\[y=mx+b\]

OpenStudy (ivettef365):

Find the slope an substitute for m, you already have b = -2

OpenStudy (ajprincess):

Let the equation be y=mx+c \[m=\frac{y_2-y_1}{x_2-x_1}\] After finding the value of m, choose one of the point and plug in the values of it and m in y=mx+c to find c. Does that help? @iribeiro1602

OpenStudy (anonymous):

yes, sort of... but now you're finding the equation :) so find the slope first! like how we did it last time :)

OpenStudy (anonymous):

...OK just a sec...

OpenStudy (anonymous):

so after you find the slope, you can narrow your choices down a bit :) and y=mx+b, the b part is the y-intercept, so afterwards, you just do what @ajprincess said, and find it and plug in! :) by substituting :)

OpenStudy (anonymous):

I've got y= -0.5x - 2

OpenStudy (anonymous):

That = y= -1/2x - 2

OpenStudy (ajprincess):

yup right:)

OpenStudy (anonymous):

yup! that's right! :) cuz 0-(-2) / -4/0 = 2/-4= -1/2 which is your slope :) and of those two values, the y values are 0 and -2, and of your choices, -2 is it! so good job :)

OpenStudy (ajprincess):

well done:)

OpenStudy (anonymous):

Thanks both!!!!!!!

OpenStudy (ajprincess):

Welcome:)